OpenStudy (anonymous):

Can help me please? Evaluate. 2 – 4 • 3 + 6 • 3

TheSmartOne (thesmartone):

Do you know PEMDAS? :)

OpenStudy (anonymous):

um isn't it like, Parenthese, Exponents, Multiplication, Division, Addition, Subtraction?

TheSmartOne (thesmartone):

yes :)

TheSmartOne (thesmartone):

so first we need to do multiplication :)

TheSmartOne (thesmartone):

from left to right :)

OpenStudy (anonymous):

okay! so mulitply 4 by 3 and then 6 by 3?

OpenStudy (anonymous):

12 and 18

TheSmartOne (thesmartone):

ok so now we have 2 - 12 + 18

TheSmartOne (thesmartone):

now we do addition and subtraction from left to right :P

OpenStudy (anonymous):

2-12= -10 -10 + 18= 8

OpenStudy (anonymous):

is that right?

TheSmartOne (thesmartone):

correct! :D

OpenStudy (anonymous):

yay! thanks

TheSmartOne (thesmartone):

Anytime! :D